Please Help:
Find the proceeds. Assume 365 days in a year. Round to the nearest cent.
$5000; discount rate 9%;
length of loan 3 months
A) $112.50
B) $4550.00
C) $4887.50
D) $4925.00

Answer by rfer(16322)

5000*.09=450 interest per yr
three months interest is 450/4=112.50
5000-112.50= 4887.50 (C)